# Isometries of spacetimes without observer horizons

### v1

@@ 4. Further results
In this section, we prove a result on uniform Lipschitz continuity of isometries, and give an application thereof to spacetimes with regular cosmological time function (where we do not require the NOH).
@@ 4.1. Uniform Lipschitzianity of isometries
where the domain of definition is $[0,\infty)$ because $\|V_{i}\|_{h}\to\infty$. We reach a contradiction to non-total imprisonment, because then $\sigma$ would be an inextendible causal curve contained in the compact set $K^{\prime}$.
@@ 4.2. Spacetimes with a regular cosmological time function
Let $\mathrm{Co}^{1,n}$ be the (solid) lightcone of the Minkowski spacetime $\mathbb{R}^{1,n}$ (i.e., the chronological future $I^{+}(0)$ of 00). It is foliated by homothetics of hyperbolic space $\mathbb{H}^{n}$ (see Figure 4). Let $\Gamma$ be a co-compact (torsion-free) lattice in the Lorentz group $\mathrm{SO}(1,n)$, meaning that $\mathbb{H}^{n}/\Gamma$ is a closed hyperbolic manifold. The quotient $M_{\Gamma}=\mathrm{Co}^{1,n}/\Gamma$ is globally hyperbolic with compact Cauchy surfaces. Furthermore, it is flat, i.e., locally isometric to Minkowski. In fact, $M_{\Gamma}$ is a warped product $M_{\Gamma}=\mathbb{R}^{+}\times_{w}S$, with $S=\mathbb{H}^{n}/\Gamma$ and warping function $w(r)=r^{2}$ (as in the case of polar coordinates in Euclidean spaces). The radius function (i.e., the projection onto $\mathbb{R}^{+}$) turns out to be the cosmological time of $M_{\Gamma}$. Finally, $M_{\Gamma}$ satisfies the NOH.

### v2

@@ 4. Further results
In this section, we prove a result on uniform Lipschitz continuity of isometries, and give an application thereof to spacetimes with regular cosmological time function (where we do not require the NFOH).
@@ 4.1. Uniform Lipschitzianity of isometries
where the domain of definition is $0,\infty)$ because $\|V_{i}\|_{h}\to\infty$. We reach a contradiction to non-total imprisonment, because then $\sigma$ would be an inextendible causal curve contained in the compact set $K^{\prime}$.
@@ 4.2. Spacetimes with a regular cosmological time function
Let $\mathrm{Co}^{1,n}$ be the (solid) lightcone of the Minkowski spacetime $\mathbb{R}^{1,n}$ (i.e., the chronological future $I^{+}(0)$ of 0). It is foliated by homothetics of hyperbolic space $\mathbb{H}^{n}$ (see Figure 4). Let $\Gamma$ be a co-compact (torsion-free) lattice in the Lorentz group $\mathrm{SO}(1,n)$, meaning that $\mathbb{H}^{n}/\Gamma$ is a closed hyperbolic manifold. The quotient $M_{\Gamma}=\mathrm{Co}^{1,n}/\Gamma$ is globally hyperbolic with compact Cauchy surfaces. Furthermore, it is flat, i.e., locally isometric to Minkowski. In fact, $M_{\Gamma}$ is a warped product $M_{\Gamma}=\mathbb{R}^{+}\times_{w}S$, with $S=\mathbb{H}^{n}/\Gamma$ and warping function $w(r)=r^{2}$ (as in the case of polar coordinates in Euclidean spaces). The radius function (i.e., the projection onto $\mathbb{R}^{+}$) turns out to be the cosmological time of $M_{\Gamma}$. Finally, $M_{\Gamma}$ satisfies the NOH.